(D^ 2 -40D' +4D'^2)Z=e^ (2x+y)

1
Expert's answer
2022-01-31T15:28:29-0500

"(D^2-4DD'+4D'^2)z=e^{2x+y}\\
\text{The auxilliary equation is }\\
m^2-4m+4=0\\
(m-2)^2=0\\
m=2 \text{ twice}\\
\text{The complementary function is:}\\"


"f_1(y+2x)+x \cdot f_2(y+2x)"


To get the particular integral.

"P.I=\frac{1}{(D^2-4DD'+4D'^2)}e^{2x+y}\\
P.I= \frac{1}{(D-2D')^2}e^{2x+y}\\
P.I=\frac{x^2}{1^2\cdot \lfloor{2}}e^{2x+y}\\
P.I=\frac{x^2}{2}e^{2x+y}"

Hence, the complete solution is the sum of the complementary function and the particular integral.


"=f_1(y+2x)+x \cdot f_2(y+2x)+\frac{x^2}{2}e^{2x+y}"